OpenStudy (anonymous):

What do all these equal up? you minus 1.)$40,520 -$35,080 2.)$37,025 -$34,492 3.)$35,175 -$32,490 4.)$42,580 -$39,687 5.)$50,032 -$45,510 6.)$48,365 -$46,045 7.)$42,115 -$38,518 8.)$39,058 -$36,427 9.)$40,185 -$37,963 10.)$38,813 -$35,148 11.)$42,561 -$38,306 12.)$45,064 -$41,570

OpenStudy (anonymous):

@allie_bear22 @AshleyWagnerCA @kelliegirl33 @ryan123345 @So1D @pooja195

OpenStudy (anonymous):

I need to check my answers it's for a portfolio okay

OpenStudy (anonymous):

Give me a minute....

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay i can try to help

OpenStudy (anonymous):

okay @AshleyWagnerCA what grade are you in

OpenStudy (anonymous):

what subject in math is it? 9th

OpenStudy (anonymous):

oh and you go to CA

OpenStudy (anonymous):

1. 40520 - 35080 = 5440 2. 37025 - 34492 = 2533 3. 35175 - 32490 = 2685 4. 42580 - 39687 = 2893 5. 50032 - 45510 = 4522 6. 48365 - 46045 = 2320 7. 42115 - 38518 = 3597 8. 39058 - 36427 = 2631 you want to help @AshleyWagnerCA .....can you do the last 4 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what state

OpenStudy (anonymous):

9. 40185 - 37963=2222 10. 38813 - 35148=3665 11. 42561- 38306=4255 12. 45064 - 41570=3494

OpenStudy (anonymous):

there you go @heyheyhelpme

OpenStudy (anonymous):

I got the same on all to okay now i need a line graph

OpenStudy (anonymous):

I am no good at graphs...sorry :(

OpenStudy (anonymous):

it's okay @kelliegirl33

OpenStudy (anonymous):

thank you for your help thoe

OpenStudy (anonymous):

i can help with that. Oh and @heyheyhelpme i go to MTSMCA Connections Academy in Minnesota

OpenStudy (anonymous):

what does MTSMCA stand for ?

OpenStudy (anonymous):

Basically, don't you just make a number graph from numbers 1000 - 6000. Then plot the points (numbers) on there. I do not know if they have to be in chronological order or numerical order.

OpenStudy (anonymous):

@AshleyWagnerCA knows because we go to the same school

OpenStudy (nathan917):

Hello guys!.

OpenStudy (anonymous):

hey @heyheyhelpme

OpenStudy (nathan917):

http://www.mathgoodies.com/lessons/graphs/images/line_example1.jpg So you want the number to all fit into a line graph if so ok.

OpenStudy (anonymous):

yeah but I need i need a line graph of all the number at the end with the >< things okay please <:~) 1. 40520 - 35080 = >5440< 2. 37025 - 34492 = >2533< 3. 35175 - 32490 = >2685< 4. 42580 - 39687 = >2893< 5. 50032 - 45510 => 4522< 6. 48365 - 46045 = >2320< 7. 42115 - 38518 = >3597< 8. 39058 - 36427 = >2631< 9. 40185 - 37963=>2222< 10. 38813 - 35148=>3665< 11. 42561- 38306=>4255< 12. 45064- 41570=>3494<

OpenStudy (anonymous):

@nathan917 can you show my how to do a line graph on microsoft word

OpenStudy (nathan917):

Im kinda stuck on this. As well as ash was.